OpenStudy (anonymous):

What is the resulting ordered pair if the value of the independent variable is 2? f(x)= -5x 4

OpenStudy (mathstudent55):

Just substitute x with 2 in -5x + 4 x = 2 y = f(2) = -5(2) + 4 (evaluate this to get y